3. Prove that (f = 0()] ^ [9 = 0(h)) = f =0(h).
Answer #1

3. since f = 0(g) → 1 f600] > Kig) tayo g = 000) = 19(80) = g(x) = Kzhe) Vands by defo of of Big 0 g(x) (3C) e b (2) must be